OpenStudy (anonymous):

can someone please explain how you would find the exact value of 3 --- √8

OpenStudy (anonymous):

it is exact you can rewrite in simplest radical form by multiplying top and bottom by \(\sqrt{2}\)

OpenStudy (anonymous):

the decimal of that is 1.414213562373095

OpenStudy (anonymous):

you get \[\frac{3}{\sqrt{8}}\times \frac{\sqrt{2}}{\sqrt{2}}=\frac{3\sqrt{2}}{\sqrt{16}}=\frac{3\sqrt{2}}{4}\]

OpenStudy (anonymous):

both numbers are "exact" second one is in simplest radical form. decimal is an approximation, not exact

OpenStudy (anonymous):

thanks that helped a lot! :D